OpenStudy (anonymous):

Solve the given linear system 5x+10y=70 5x+25z=270 10y+25z=300

OpenStudy (hitaro9):

There are a few ways you can solve this, either by substituion or by straight addition/substraction. I always thought substitutionw as easier to understand, so I'll go with that. You can simplify the 2nd equation to be x= 54-5z, I assume you can do the algebra there.

OpenStudy (hitaro9):

If you take that, and apply the same principle to a few others, you can get values for y and z as well. for example, the first equation you can get x= 14-2y

OpenStudy (hitaro9):

Just do that for a couple of values, solving for x,y,z and z, and then resubstitue htose values back in to different equations when necesary.

OpenStudy (hitaro9):

Does that make sense?

OpenStudy (hitaro9):

Cause you're taking some value, finding what it is in other's terms, and just replugging it in

OpenStudy (anonymous):

woah..

OpenStudy (anonymous):

very confusing at all.. :(

OpenStudy (hitaro9):

What're you finding confusing? Go ahead and try rearranging some of the equations to get some simpler equations such as y=7-1/2x

OpenStudy (hitaro9):

That's the first step, should be a simple step, basic algebra getting the variable on one side

OpenStudy (anonymous):

uh.. ive never try like this.. i only tried like this. y=5x-3 y=3x-1

OpenStudy (hitaro9):

Okay, using that example, you know that y= both of those things right? So they have to equal eachother. Cause if y=n and y=m, then n must = m, right? So if you set those equal to each other 5x-3=3x-1 Then just simplify 2x = 2 x =1 You just need to do that, but with bigger numbers and a z thrown in

OpenStudy (hitaro9):

Does that make sense or no?

OpenStudy (anonymous):

but.. what if there are three? could you give me an example?

OpenStudy (hitaro9):

It's just a lot more tedious work. If you solve both y and z for x, the plug them in to the third equation, then you can get x For example, if y= 2x and z= 2x and y+z =4 Then you can plug those values in 2x +2x =4 4x=4 x= 1 And then you can use the fact that x =1 to find the original 2, so if y=2x then y=2.

OpenStudy (hitaro9):

You can do the same with your problem, it just looks like its going to have fractions and/or big numbers and I don't have a calculator or pencil + paper handy, but just follow the same steps

OpenStudy (hitaro9):

Just solve y and z for x, plug whatever those 2 equations are in to the 3rd, solve for x, and then use that value of x to find the original 2.

OpenStudy (anonymous):

thanks :)

OpenStudy (hitaro9):

No problem. Let me know if you need any more help. You can post what you did and I can double check that you did everything right. This is just how I do those problems by the way. There are other methods that your teacher may prefer, this is just the one that I think makes the most sense.